OpenStudy (anonymous):

simplify this expression (3-4i) (5+2i)

OpenStudy (anonymous):

23-14i

OpenStudy (anonymous):

are u for sure

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

like step by step

OpenStudy (anonymous):

@Sumanth4phy is correct

OpenStudy (anonymous):

ok i beleive he is i need step by step though

OpenStudy (anonymous):

Multiply them: 15+6i-20i-8i^2------->i^2=-1 15-14i-8(-1) 15-14i+8=23-14i

OpenStudy (anonymous):

\[(3-4i)(5+2i) = 15+6i-20i-8i^2\] Since we know that \[i^2=\sqrt{-1}^2=-1\]The expression gives \[15-14i+8=23-14i\]

OpenStudy (anonymous):

thank you you guys are awesome

OpenStudy (anonymous):

You are wlcm!